OpenStudy (anonymous):

The third term of an arithmetic sequence is 9 and the seventh term is 31. Find the sum of the first twenty-two terms.

OpenStudy (helder_edwin):

an arithmetic sequence is of the form \[ \large a_1\qquad\text{not given} \] \[ \large a_2=a_1+d \] \[ \large 9=a_3=a_2+d=a_1+2d \] so \[ \large 31=a_7=a_1+6d \]

OpenStudy (helder_edwin):

u end up with \[ \large 9=a_1+2d \] \[ \large 31=a_1+6d \] solve this system

OpenStudy (helder_edwin):

r u there @katiebugg ?

OpenStudy (anonymous):

sorry yeahh was working on a prob thanks!!!